I have to determine whether the graphs of the equation are perpendicular, parallel, or neither. 4x+4y=-4 and 4x-4y=0. I put perpendicular, is this correct?
Answer by tanimachatterjee(60)

equations are 4x+4y =-4.....................(1)
and 4x-4y=0...........(2)
so we can write the equation (1) as 4y = -4x -4
or y = -x -1
so the slope is -1
(2) we can write as 4y = 4x
or y=x
slope =1
product of these 2 slope is (-1)(1) = -1
so they r perpendicular to each other
and your ans is ABSOLUTELY correct